OpenStudy (anonymous):

Solve 2x^2=14, I got sqrt(7), -sqrt(7). Now it says, (and the part I don't get is "Find the x-intercepts of f(x)2x^2-14"

OpenStudy (anonymous):

This question has 2 parts.

OpenStudy (anonymous):

its the same thing

OpenStudy (agreene):

yep, it's the same.

OpenStudy (anonymous):

its just 2x^2-14=0

OpenStudy (agreene):

since the first step in finding of the roots takes you to: 2x^2=14

OpenStudy (anonymous):

I'm using mathlab, so do I just type in the same thing?

OpenStudy (anonymous):

For the x-intercepts...

OpenStudy (anonymous):

yes i think so

OpenStudy (agreene):

x intercepts are: \[\pm \sqrt 7\] enter that however you need to :D

OpenStudy (anonymous):

Ah! :o What does the first symbol mean? Two of sqrt(7)? I hope this works.

OpenStudy (anonymous):

it mean plus or minus